Razor:

A 20,000 gallon swimming pool contains 5000 gallons of water when a hose is placed in the pool and begins adding water at a rate of 1250 gallons per hour.

Razor:

1 attachment
jhonyy9:

Razor any idea ? you need creat a function with variable x what f(x) will start from 5000 do you understand me ?

Razor:

No I don't understand any of it I missed the class for this lesson....

jhonyy9:

you know that f(0) = 5000 gallon and after every hours the pool get 1250 gallon water so after the first hour the pool contain 5000 +1250 gallon water = 6250 gallon after second hour the pool contain 6250+1250 = 7500 gallon water do you understand me ?

Razor:

ohhhh....so I just keep adding 1250 because its like ...per hour?

jhonyy9:

@ThisGirlPretty yes exactly

jhonyy9:

and using this you ned make a diagramm

jhonyy9:

need

jhonyy9:

it's ok now ?

Razor:

erm....do we use rise and run?

jhonyy9:

@ThisGirlPretty can you please help in explain it ?

Razor:

lolz.....sorry dewd

jhonyy9:

the diagramm start from point f(0) = 5000 so y = 5000 after 1 hour so x=1 the y=5000+1250 = 6250 and so much continue

jhonyy9:

when x=2 so y = 7500

Razor:

..............................................................maybe I should use khan academy o-o or wait for dude o-o

jhonyy9:

why ? just us what i ve wrote above and drawn the diagramm till the pool will be full arive at 20000 gallon of water

jhonyy9:

@ThisGirlPretty please can you explaining more understandably ?

Razor:

one sec I'm gonna see if I can do it and show you

jhonyy9:

Razor this diagramm is very very easy just us what i ve wrote above

Razor:

I'm more of a visual learner. Sorry. Um is this right?

1 attachment
Razor:

Or do you mean diagram like a shape

jhonyy9:

no just a line a curve

Razor:

ohhh

Razor:

Yeah I'm still cofused and my line is all over the place

1 attachment
Razor:

I def didn't do it right

dude:

Oh, its not exponential, its linear

Razor:

o-o ...but howz I dew dis

dude:

So jhonny said that when x is 0, the y is 5,000 So 5,000 is your starting points

jhonyy9:

exactly dude - thank you thet help in explaination

Razor:

ohhhhh

Razor:

okie okie

dude:

`1250 gallons per hour.` This is your rate (rise/run) per is like saying over (a fraction) So, 1250 \(\color{red}{per}\) hour =>\(\frac{1250}{hour}=>\frac{1250}{1}\)

Razor:

..... I'm guessing I'm going to stop my line at 4 hours... and then continue going up from there?

dude:

Add 1250 each time to get your next point 1542234494-5beca17e2b18d75fe8d38c-image.png That is a rough estimate, I only plotted the first hour

dude:

Starting = 5,000 Point 1 = 5,000+1250 = 6250 Point 2 = 6250 + 1250 = 7500 and so on...

Razor:

oh so I literally start ON the line..

dude:

Yes

Razor:

1 attachment
Razor:

Like that and then so on?

dude:

to to the next point, just add 1250

dude:

When re-check your point

dude:

Unless you are actually at 6250

Razor:

It won't let me go inbetween the numbers lol is this right?

1 attachment
dude:

Oh, hmm can you plot points on 3 different locations? Or does it have to be every x value has a point?

Razor:

Like this you mean?

1 attachment
dude:

No no, I mean like only have a point for example when x is 2 1542235331-5beca4c33875b829f0ec9000-image.png

Razor:

ohhh no it doesn't let me do that

dude:

Welp, I guess the way you were doing it is right though it isn't technically right Every point is supposed to go up by 1250, not 1000

Razor:

lolz ... okie I can't make it go that little or far though

dude:

You're only supposed to go up to when y is 20,000, since that is the capacity of the gallon

Razor:

1 attachment
dude:

Up to 20,000 Also its a straight line, there are no jumps

Razor:

1 attachment
Razor:

Ah I got it

dude:

Oh wait you can play specific points

Razor:

??

dude:

You want to post these main ones 1542235904-5beca7002dd317c96d4d9a00-image.png

dude:

Oh it only goes up by 10,000 Only plot (4, 10000) and (12, 20000)

Razor:

ah okie XD

dude:

Yes

Razor:

errrrr points all look wonky

dude:

How so?

Razor:

1 attachment
dude:

You don't need the other line The points in your answer should only be (0, 5000) (4, 10000) (12, 20000)

Razor:

1 attachment
Razor:

Took way longer than I thought

dude:

Perf

dude:

Yes sorry

Razor:

XD iz fine my fault

dude:

D: Well that is your graph

Razor:

okiez.. do yew think yew could stay on for more math questions?

dude:

Sure

Razor:

okie thanks